A hydrogen atom be made up of a single proton and a single electron.the proton has a charge of+e and electron has +e.in the ground state of the atom the electron orbits the proton at most probable distance of 5.39*10^-11m determine the electric force on the electron due to the proton
1
Expert's answer
2020-05-12T09:59:56-0400

Let us determine the electric force between two particles with charges "q_1" and "q_2" . According to the Coulomb's law (see https://en.wikipedia.org/wiki/Coulomb%27s_law) force can be calculated as

"F = k_e\dfrac{q_1q_2}{r^2}," where "k_e" is Coulomb's constant ("k_e = 9\times10^9 \,N\cdot\,m^2/C^2" ) and "r" is the distance between charges.

Therefore, we may calculate the module of force

"F = 9\times10^9\dfrac{(1.60\cdot10^{-19})\cdot(1.60\cdot10^{-19})}{(5.39\cdot10^{-11})^2} = 7.93\cdot10^{-8}\,N."